### v2

@@ Abstract
With the advancement of multi-robot technology, cooperative exploration tasks have garnered increasing attention. This paper presents a comprehensive review of multi-robot cooperative exploration systems. First, we review the evolution of robotic exploration and introduce a modular research framework tailored for multi-robot cooperative exploration. Based on this framework, we systematically categorize and summarize key system components. As a foundational module for multi-robot exploration, the localization and mapping module is primarily introduced by focusing on global and relative pose estimation, as well as multi-robot map merging techniques. The cooperative motion module is further divided into learning-based approaches and multi-stage planning, with the latter encompassing target generation, task allocation, and motion planning strategies. Given the communication constraints of real-world environments, we also analyze the communication module, emphasizing how robots exchange information within local communication ranges and under limited transmission capabilities. Finally, we discuss the challenges and future research directions for multi-robot cooperative exploration in light of real-world trends. This review aims to serve as a valuable reference for researchers and practitioners in the field.
@@ 1 INTRODUCTION
Initially, multi-robot cooperative exploration systems are limited to handling lab settings or very simple environments, such as small offices [11]. Driven by real-world needs, these systems have evolved to handle large-scale and extreme outdoor environments, such as caves [1, 2, 3], underground tunnels [4], forests [5], and playgrounds [6] (shown in Fig. 1).
To advance the development of multi-robot cooperative exploration systems for large-scale and extreme outdoor environments, the DARPA Subterranean (SubT) Challenge [12, 13] has emerged. The complexity of the DARPA SubT Challenge originates from four core challenges: denied GNSS signals, large-scale scenarios, tough terrains, and intermittent communication. These challenges drive the development of heterogeneous robot teams to improve system robustness and exploration efficiency, which also becomes a challenge. All the challenges guide the researchers to develop a multi-robot cooperative exploration system with accurate localization and map fusion, efficient cooperative mechanism, representation methods with small data transmission volume and high precision, and efficient communication strategies. In the DARPA SubT Challenge, various research teams have developed and deployed innovative approaches for multi-robot cooperative exploration. Notable contributions include Team CERBERUS’s CompSLAM architecture [14], Team CoSTAR’s Nebula architecture [15], Team CSIRO’s Wildcat architecture [16], and the system developed by Team CTU-CRAS-NORLAB [17], along with other significant works [18, 19, 20]. These efforts have introduced robust and practical solutions, advancing the design and implementation of multi-robot exploration systems for extreme and large-scale environments.
Developing multi-robot cooperative exploration systems for extreme and large-scale environments presents several challenges, which can be broadly categorized as follows:
@@
Based on the summary of the main challenges faced by multi-robot cooperative exploration systems when exploring extreme and large-scale outdoor environments, we divide the technical composition of multi-robot cooperative exploration systems into three corresponding modules: localization and mapping, cooperative planning, and communication. Each module is designed to address the specific difficulties. Fig. 2 describes the coordination between the modules of the multi-robot cooperative exploration system. First, the robots need to perform self-localization and determine their relative positions through localization techniques. Next, they construct an environmental map based on sensor data, which serves the subsequent path planning and navigation. On this basis, the robots perform cooperative planning by considering the current state, environmental information, and other factors, to explore the unknown environment efficiently. Throughout the process, the communication module facilitates information exchange, such as transmitting pose data and local maps. These modules are interrelated and mutually influence each other, forming the technical framework for multi-robot cooperative exploration.
@@
The structure of the paper is as follows: Section 2 provides an overview of the technical background of multi-robot cooperative exploration systems, reviews their development history, and discusses the related tasks in detail. Section 3 introduces localization and map fusion methods used to extend the map boundaries in multi-robot cooperative exploration systems. Section 4 presents the cooperative planning methods for multi-robot systems, including multi-stage planning and learning-based planning. Section 5 discusses the robot interaction issues under limited communication range and bandwidth conditions. Section 6 summarizes the current unresolved issues and explores potential future research directions. Finally, Section 7 concludes the paper by summarizing the research content. Fig. 3 shows the overview structure of multi-robot cooperative exploration systems in this survey.
@@ 6 Open Problems and Possible Future Research Directions
VI-A-1. Exploring in Extreme Environments

### v3

@@ Abstract
With the real need of field exploration in large-scale and extreme outdoor environments, cooperative exploration tasks have garnered increasing attention. This paper presents a comprehensive review of multi-robot cooperative exploration systems. First, we review the evolution of robotic exploration and introduce a modular research framework tailored for multi-robot cooperative exploration. Based on this framework, we systematically categorize and summarize key system components. As a foundational module for multi-robot exploration, the localization and mapping module is primarily introduced by focusing on global and relative pose estimation, as well as multi-robot map merging techniques. The cooperative motion module is further divided into learning-based approaches and multi-stage planning, with the latter encompassing target generation, task allocation, and motion planning strategies. Given the communication constraints of real-world environments, we also analyze the communication module, emphasizing how robots exchange information within local communication ranges and under limited transmission capabilities. In addition, we introduce the actual application of multi-robot cooperative exploration systems in DARPA SubT Challenge. Finally, we discuss the challenges and future research directions for multi-robot cooperative exploration in light of real-world trends. This review aims to serve as a valuable reference for researchers and practitioners in the field.
@@ 1 INTRODUCTION
Initially, multi-robot cooperative exploration systems are limited to handling lab settings or very simple environments, such as small offices [11]. Driven by real-world needs, these systems have evolved to handle large-scale and extreme outdoor environments, such as caves [1, 2, 3], underground tunnels [4], forests [5], and playgrounds [6] (shown in Fig. 1). The complexity of these environments originates from four core challenges: denied GNSS signals, large-scale scenarios, tough terrains, and intermittent communication. These challenges drive the development of heterogeneous robot teams to improve system robustness and exploration efficiency, which also becomes a challenge. Developing multi-robot cooperative exploration systems for extreme and large-scale field environments presents several challenges, which can be broadly categorized as follows:
@@
All the challenges guide the researchers to develop a multi-robot cooperative exploration system with accurate localization and map fusion, efficient cooperative mechanism, representation methods with small data transmission volume and high precision, and efficient communication strategies. Based on the summary of the main challenges faced by multi-robot cooperative exploration systems when exploring extreme and large-scale outdoor environments, we divide the technical composition of multi-robot cooperative exploration systems into three corresponding modules: localization and mapping, cooperative planning, and communication. Each module is designed to address the specific difficulties. Fig. 2 describes the coordination between the modules of the multi-robot cooperative exploration system. First, the robots need to perform self-localization and determine their relative positions through localization techniques. Next, they construct an environmental map based on sensor data, which serves the subsequent path planning and navigation. On this basis, the robots perform cooperative planning by considering the current state, environmental information, and other factors, to explore the unknown environment efficiently. Throughout the process, the communication module facilitates information exchange, such as transmitting pose data and local maps. These modules are interrelated and mutually influence each other, forming the technical framework for multi-robot cooperative exploration.
@@
The structure of the paper is as follows: Section 2 provides an overview of the technical background of multi-robot cooperative exploration systems, reviews their development history, and discusses the related tasks in detail. Section 3 introduces localization and map fusion methods used to extend the map boundaries in multi-robot cooperative exploration systems. Section 4 presents the cooperative planning methods for multi-robot systems, including multi-stage planning and learning-based planning. Section 5 discusses the robot interaction issues under limited communication range and bandwidth conditions. Section 6 introduces the actual application of multi-robot cooperative exploration systems in field environments. Section 7 summarizes the current unresolved issues and explores potential future research directions. Finally, Section 8 concludes the paper by summarizing the research content. Fig. 3 shows the overview structure of multi-robot cooperative exploration systems in this survey.
@@ V-B-2. Adaptive State-aware Strategy
6 The Field Application of Multi-robot Cooperative Exploration Systems
Unlike laboratory research that seeks innovation and superior performance, field multi-robot cooperative exploration systems emphasize robustness, stability, and efficiency. Indoor environments have sparse obstacles, so single-type robot exploration systems can explore them effectively. However, in extreme and complex field settings like underground mines, tunnels, or underwater areas, challenges such as limited visibility, communication constraints, rough terrain, and GNSS-denied make single-type robot systems inefficient. This drives the need for heterogeneous robot systems. heterogeneous robots have different motion models, dynamics, payload capacities, computational resources, and endurance. Therefore, tasks must be allocated according to each robot’s strengths to enhance exploration efficiency.
This section selects the work of excellent teams in the US Defense Advanced Research Projects Agency (DARPA) Subterranean (SubT) Challenge [221, 222], including MARBLE [223], Explorer [224], CTU-CRAS-NORLAB [225], CSIRO Data61 [226], CoSTAR [227], and CERBERUS [228], and discusses the design of heterogeneous multi-robot cooperative exploration systems and their applications for exploring extreme, complex, and large-scale field environments.
VI-A. Heterogeneous Robots in Multi-robot Cooperative Exploration Systems
A heterogeneous multi-robot system is a team of robots that differ in type, with differences in shape, mobility, and sensor configuration [229]. This diversity enables the system to perform tasks efficiently in complex, unknown, and dangerous field environments through complementary advantages. In the DARPA SubT Challenge, many teams demonstrated how to use heterogeneous multi-robot cooperative exploration systems to deal with various challenges in underground environments. Based on the mobility of robots, these systems mainly consist of three types: UGV (unmanned ground vehicles), legged, and UAVs. UGV can be further divided into wheeled and tracked categories.
Wheeled robots usually have high speed and long endurance, making them ideal for moving quickly on flat or slightly rough terrain. Their higher payload capacity allows them to carry more sensors and equipments. They can rapidly cover large areas to obtain environmental information and deploy communication nodes to enhance the network. This makes them well-suited for perception, mapping, and communication node deployment tasks. Examples include the Husky robot used by CoSTAR and CTU-CRAS-NORLAB, Armadillo from CERBERUS, and R1/R2 from Explorer.
When moving over very rough terrain, wheeled robots often struggle due to their limited obstacle - crossing ability, which has led to the need for tracked robots. Tracked robots can handle rough, muddy, or soft ground well. They are stable, can carry many sensors and devices, and are suitable for long - term tasks in complex terrain, such as perception, mapping, and communication node deployment. Examples are the Absolem from CTU - CRAS - NORLAB, the BIA5 ATR from CSIRO Data61, and the SuperDroid LT2 - F. However, since tracked robots are generally slower than wheeled ones, it’s important to analyze and decide which robot to use for exploring different terrains to ensure efficient exploration.
Legged robots like quadrupedal and hexapodal robots can stably walk on stairs, slopes, and uneven or muddy terrain due to their strong adaptability. Their flexibility and mobility make them suitable for complex and narrow areas. With a certain payload capacity, they can carry multiple sensors and computing units, enabling highly autonomous navigation and exploration. Examples include CTU-CRAS-NORLAB’s PhantomX Mark II,CSIRO Data61’s Hexapod, Ghost Robotics’ Vision60, and CERBERUS’s ANYmal B300.
UAVs can move quickly in the air, unrestricted by ground obstacles. Equipped with light sensors like cameras and LiDAR, they offer an aerial perspective. They help the team understand the overall environment, assist other robots in path planning, and explore narrow or dangerous areas such as shafts and crevices. UAVs can also act as communication relay nodes to expand network coverage. To address their limited endurance, they can be transported on wheeled or legged robots and deployed when needed, as demonstrated by Explorer’s DS.
The multi-robot cooperative exploration system with heterogeneous robots can combine the strengths of different robot types during exploration, reduce the weaknesses of single-type robots, and enhance system stability, robustness, and exploration efficiency. However, it also causes problems like ensuring consistent localization estimation, aligning multi-modal perception and maps, planning multi-agent collaboration, and designing a reliable communication network. Existing possible solutions to these problems have been described in Section 3,  4,  5, so won’t repeat here.
VI-B. The Field Application of Multi-robot Cooperative Exploration Systems
In the DARPA Subterranean Challenge, heterogeneous multi-robot cooperative exploration systems have unique advantages in dealing with complex and variable outdoor environments such as tunnels, urban, and caves because of the limits of these environments. In such environments, balancing robustness, stability, and efficiency of multi-robot cooperative exploration systems is crucial because the environments are full of uncertainties and challenges. Over-relying on a useful but complex module can limit system performance due to the buckets effect. Therefore, each module should be simple and practical while ensuring close cooperation. In addition, tasks should be allocated according to the characteristics of heterogeneous robots to give full use to their strengths.
The CoSTAR framework’s Networked Belief-aware Perceptual Autonomy (NeBula) system comprises the following modules: a localization and mapping module consisting of Heterogeneous and Resilient Odometry estimator (HeRO) [230] and Large-scale Autonomous Mapping and Positioning (LAMP) [231], a cooperative planning module Probabilistic Local and Global Reasoning on Information roadMaps(PLGRIM) [232], and a communication module Collaborative High-bandwidth Operations with Radio Droppables (CHORD) [233]. HeRO enhances localization resilience through parallel processing of multiple sensors and algorithms, such as integratig LiDAR-inertial (LIO), visual-inertial (VIO), thermal-inertial (TIO) kinematic-inertial (KIO), contact-inertial (CIO) and RaDAR-inertial (RIO), while considering redundancy and heterogeneity. This architecture enables fault detection and system adaptation to ensure accurate and stable positioning in GNSS-denied field environments. LAMP serves as a factor graph-based SLAM solution designed for large-scale, perceptually degraded environments, such as those encountered in the DARPA Subterranean Challenge. It achieves consistent global environmental representation by fusing multi-sensor inputs to detect loop closures and mitigate accumulated errors. PLGRIM employs a hierarchical Information RoadMap (IRM) representation and planning architecture to enable efficient exploration in expansive environments. The system models autonomous exploration as a partially observable Markov decision process (POMDP), enabling uncertainty-aware decision-making. By integrating local and global planning strategies, PLGRIM dynamically adjusts robotic exploration paths to maximize information gain while minimizing action costs. CHORD maintains high-bandwidth communication links among multiple robots, ensuring effective command execution, autonomous operations, and data collection in communication-constrained field environments. The module extends wireless mesh networks through deployable communication nodes, enabling extended-range connectivity between robots and base stations. Besides, NeBula implements mission-specific role allocation based on robotic platform capabilities and environmental characteristics: tracked and hybrid-powered robots navigate complex mine and tunnel terrains, while UAVs perform rapid data relay or vertical shaft exploration. This heterogeneous coordination leverages each platform’s unique advantages to optimize system performance.
In the localization and mapping module, the CERBERUS system employs a complementary hierarchical multi-modal sensor fusion approach (CompSLAM) [234], integrating data from visual/thermal cameras, LiDAR, and inertial sensors to achieve robust robot pose estimation. This fusion strategy ensures reliable localization even under degraded perceptual conditions like geometric self-similarity, darkness, or obscurants. Additionally, CERBERUS adopts a centralized multi-robot multi-modal mapping (M3RM) [235] optimization method, where submaps from individual robots are aggregated and refined at the base station using factor graphs with visual, LiDAR, and prior constraints. This approach not only enhances the localization accuracy of individual robots but also improves the system’s environmental perception consistency, enabling artifact position refinement in the DARPA coordinate frame. For the cooperative planning module, CERBERUS uses Graph-based exploration path planner (GBPlanner) [236] - a unified graph-search-based exploration path planning method applicable to both legged and aerial robots. This method guides robots to autonomously explore complex environments efficiently through hierarchical local-global planning with volumetric information gain metrics. During planning, the system generates optimal exploration paths based on occupancy maps, traversability constraints from elevation maps, and frontier vertices from a global graph. The planning framework also incorporates geofencing for obstacle avoidance and accounts for robot motion constraints to ensure safe navigation. Furthermore, CERBERUS implements automated homing based on remaining mission time and supports multi-robot coordination through frontier sharing in the global graph. The communication module of CERBERUS employs multiple strategies to maintain reliable connectivity between robots and the base station. For underground environments, wheeled robots equipped with high-gain antennas and fiber-optic cables establish stable communication links with the base station. Meanwhile, legged robots dynamically deploy wireless relay nodes at critical locations to extend the communication network’s coverage. This creates an adaptive, dynamic communication network during exploration missions. CERBERUS also optimizes task allocation based on robot capabilities and environmental demands. The ANYmal quadrupedal robot serves as the primary exploration platform due to its outstanding endurance and complex terrain traversal capabilities, enabling stable navigation in rugged underground environments. The Aerial Scouts multi-rotor UAVs leverage their rapid flight agility to explore confined spaces like narrow passages or vertical shafts and perform fast reconnaissance. The ArmadiIIo wheeled robot specializes in communication node deployment and object detection, providing network support via its high-gain antennas and fiber-optic cables while contributing to environmental sensing tasks.
In the localization and mapping module, CSIRO Data61 developed the Wildcat SLAM system [237], which enables each robot to generate a shared global map and achieve efficient environmental mapping and localization through multi-sensor data fusion. The Wildcat SLAM system generates SLAM frames and shares them across robots, allowing each robot to construct a globally consistent map based on both its own data and inputs from other agents. While absolute consistency across all robots’ global maps cannot be guaranteed in complex field environments, the overlapping of frames from a shared initial region enables each robot to incrementally build its global solution using all available data. For the cooperative planning module, CSIRO Data61 designed a Multi-Robot Task Allocation (MRTA) method. This approach employs a market auction mechanism, where robots bid for tasks based on expected rewards, enabling rational task distribution. Robots dynamically adjust task priorities and bidding strategies by evaluating task locations, their own states, and communication conditions with other robots. Additionally, CSIRO Data61 introduced a priority zone tool, allowing human operators to influence task allocation, thereby enhancing system flexibility and adaptability. In the communication module, CSIRO Data61 designed the Mule communication system, which utilizes a peer-to-peer network to enable robots to share data within communication range while maintaining task execution during outages. The Mule system dynamically discovers other robots in the network and establishes point-to-point connections for efficient data sharing. This strategy addresses communication challenges in subterranean environments and ensures continued collaboration post-recovery through persistent data storage and synchronization mechanisms. CSIRO Data61 assigns roles based on robotic capabilities. The BIA5 ATR robot, with its strong obstacle-crossing capability and high payload capacity, is tasked with carrying communication nodes and UAVs, while the smaller SuperDroid LT2-F robot leverages its agility and compact size to explore narrow spaces. UAVs, benefiting from their high speed and maneuverability, rapidly survey large areas or hard-to-reach zones.
The system designed by CTU-CRAS-NORLAB employs a robust Iterative Closest Point (ICP) [238] based localization and mapping method, applied across all robot types to ensure stability and reliability in complex environments. In Urban scenarios, characterized by vertically diverse structures, the team adopted the Advanced implementation of the original Lidar Odometry and Mapping [239] (A-LOAM) algorithm to process 3D LiDAR data. This method eliminates the need for prior motion estimation, making it more suitable for navigating intricate building interiors. Although challenges such as map warping and drift persist, adjustments to the algorithm and integration of auxiliary measurements reduce localization errors, thereby enhancing the robustness of the localization and mapping module. Additionally, CTU-CRAS-NORLAB developed vision-based, radio-based, and olfactory-based target localization methods, alongside adaptive traversal and information entropy-driven exploration strategies. These approaches enable robots to plan paths and explore unknown regions based on environmental data and onboard sensor inputs. By refining frontier-based exploration, the system reduces computational demands while improving environmental coverage efficiency through camera model-aware exploration waypoint generation. Multi-robot coordination is primarily achieved via 4G Mobilicom mesh transceivers, allowing operators to directly control robots or set navigation goals. Robots share positional data through low-bandwidth links to avoid redundant exploration. Furthermore, CTU-CRAS-NORLAB implemented a multi-network communication strategy combining short-range Wi-Fi, Mobilicom mid-range links, and Mote long-range links. These provide varying levels of reliability, range, and bandwidth to meet data transmission demands in challenging field environments. In Tunnel settings, larger Mobilicom modules extend communication range, while smaller modules are deployed in Urban environments, enabling all robotic platforms to act as communication nodes and enhance network coverage. Mote modules facilitate critical data sharing—such as status, location, and detection information—among robots, ensuring peer-to-peer communication even when base station connectivity is lost. CTU-CRAS-NORLAB emphasizes task allocation based on robotic capabilities: large wheeled robots excel in long-distance exploration, while compact hexapod robots navigate confined spaces. Larger robots also carry additional equipment, such as communication relays, to support the system’s communication infrastructure. This heterogeneous task distribution optimizes exploration efficiency by leveraging the unique strengths of each platform.
The multi-robot cooperative exploration system developed by Explorer ensures robustness and stability in complex field environments through modular design and rational allocation of tasks among heterogeneous robots. In the localization and mapping module, Super Odometry [240] employs an IMU as the primary sensor, integrating constraints from vision, thermal imaging, and LiDAR odometry. A selective fusion strategy automatically prioritizes effective constraints while rejecting erroneous data associations, significantly enhancing real-time performance and enabling rapid recovery from potential sensor failures. By decomposing large factor graphs into multiple sub-factor graphs such as IMU, LIO, and VIO, each sub-factor graph tightly couples raw sensor measurements internally while maintaining loose coupling between sub-factor graphs through pose constraint transmission. This architecture ensures exceptional robustness in harsh environments, allowing the system to continue operating via redundant sensor constraints even if one sensor fails. For cooperative planning, a hierarchical exploration planner manages sparse global-level information to generate coarse exploration routes, while maintaining dense local-level information to ensure comprehensive sensor coverage. The global planner operates on low-resolution maps to define exploration directions, whereas the local planner utilizes high-resolution maps for detailed path planning, achieving full coverage of localized areas. The communication module incorporates an autonomous node deployment algorithm and a ledger mechanism. The deployment algorithm evaluates communication quality between robots and deployed nodes using local point cloud data. A ray-casting method from the robot to existing nodes determines "line-of-sight visible" or "non-line-of-sight visible" status. New communication nodes are autonomously deployed when distances to existing nodes exceed a threshold, ensuring network coverage. The ledger, a distributed data structure for efficient map and object detection data sharing, records all submap metadata and broadcasts updates. Robots can request specific submaps via the ledger, which also optimizes bandwidth usage by tracking submap availability and local storage status. In heterogeneous robot task allocation, the system categorizes robots into UGVs(R1, R2, R3) and UAV(DS) based on perception and mobility capabilities. UGVs excel at long-distance, high-speed exploration and communication network deployment. In tunnel environments, R1 and R2 explore extended passages while establishing communication backbones. The UAV DS specializes in three-dimensional exploration of confined spaces. In urban settings, DS rapidly maps multi-story structures and narrow passages within buildings.
In the localization and mapping module, MARBLE utilizes tools and methodologies such as Google Cartographer [241] and Voxblox [242] to support robot positioning and mapping in unknown and complex environments. Google Cartographer, renowned for its real-time loop closure correction, ease of use, and comprehensive documentation, is employed to generate probabilistic occupancy grid representations for ground robots, enabling global path planning and situational awareness for human supervisors. Voxblox provides aerial robots with sufficient path clarity by generating a signed distance field (SDF) representation of 3D environments. For mapping, the MARBLE team leverages the open-source OctoMap package to create sparse voxel-based environmental representations and has developed a custom mapping package that incrementally integrates map segments from other robots through direct merge functionality during map creation, thereby conserving computational resources. In the cooperative planning module, MARBLE designs a metric-topological graph-based planner [243] and a continuous frontier-based planner to adapt to diverse underground environments. Additionally, MARBLE implements an internal market-based multi-agent coordination algorithm [244], enabling robots to coordinate without relying on continuous bidirectional communication. Even under message loss or communication outages, robots autonomously proceed to their lowest-cost target points as if uncoordinated. This approach mitigates system-wide efficiency degradation caused by communication issues, ensuring stability and robustness in cooperative planning. For the communication module, MARBLE integrates communication beacons, a UDP-Mesh communication network, and prioritized message handling. Communication beacons extend communication ranges and facilitate map and target point sharing among robots. The UDP-Mesh network addresses challenges of packet loss and transient connectivity in complex environments by providing discovery services and reliable User Datagram Protocol (UDP) datagram transmission. Compatible with any underlying Layer 2 mesh network solution, it requires no modifications to Robot Operating System (ROS) nodes or diagnostic tools. This module enables efficient sharing of maps, target points, and other critical data to support multi-robot cooperative planning. It also prioritizes data types, ensuring high-value data (e.g., artifact discovery reports) are transmitted first, enhancing system efficiency and robustness. MARBLE optimizes heterogeneous robot task allocation through its multi-agent coordination algorithm, which facilitates data sharing and target deconfliction. For instance, the Clearpath Husky and Superdroid HD2 ground robots, with their high payload capacity and endurance, are suited for long-duration exploration in complex terrains. In contrast, the Lumenier QAV500 aerial robot excels in rapid reconnaissance and maneuvering through confined spaces.
In addressing the numerous challenges faced by multi-robot cooperative exploration systems in comprehensive field applications, many teams have contributed unique solutions through the design of heterogeneous multi-robot cooperative exploration systems in the DARPA SubT Challenge. The technologies employed extensively reference existing mature engineering-proven solutions, thereby enhancing system stability, robustness, and operational efficiency. However, these representative works still commonly confront challenges including communication stability dependency in extreme field environments, high computational complexity, and insufficient adaptability to dynamic scenarios. Future efforts need to focus on optimizing lightweight communication protocols, enhancing computing capabilities, and exploring adaptive cooperative algorithms to balance system efficiency and robustness.
@@ 7 Open Problems and Possible Future Research Directions
VII-A-1. Exploring in Extreme Environments